OpenStudy (reganmckay32):

Country A has a growth rate of 4.2% per year. The population is currently 5,853​,000, and the land area of Country A is 11​,000,000,000 square yards. Assuming this growth rate continues and is​ exponential, after how long will there be one person for every square yard of​ land?

OpenStudy (518nad):

4.2% grown rate per year so lets say currently the population is at P, in 1 year you will have P(1+0.042) in 2 years you will have P*(1+0.042)*(1+0.042) = P*(1+0.042)^2 in n years you will have P*(1+0.042)^n

OpenStudy (518nad):

5,853​,000 (1+0.042)^n=11​,000,000,000 solve for n

OpenStudy (518nad):

Hi, do you want to continue you from here, you must use log function

OpenStudy (518nad):

remember that if u have x^n=y if you want to solve for n Start by taking log of both sides log(x^n)=log(y) now the left side simplfies to nlog(x)=log(y) n=log(y)/log(x)

OpenStudy (reganmckay32):

I see what you're saying. I'm not sure how to plug it in though.